OpenStudy (anonymous):

If the area of a triangle is 120 square centimeters, what is the length of its base?

OpenStudy (anonymous):

remember in order to find the area of a triangle u need to use the formula base times lengh divided by 2

OpenStudy (dan815):

Area = 1/2 b*h 120 = 1/2 B*h 240/h=B <---- base can be anything that satisfies this equation

OpenStudy (triciaal):

The area of a triangle is half base times height. A = 1/2 b* h b = 2A/h The question is to find b but we do not have the height, h. you have 2 unknowns

OpenStudy (anonymous):

Thank you and God Bless you both!